# Multi-stage tomography based on eigenanalysis for high-dimensional dense unitary quantum processes

### v1

@@ VI Conclusion and future works
The above approach may however be expected to yield limited performance when considering high-dimensional state spaces, due to the estimation errors of the Quantum State Tomography (QST) that it uses as a building block. We very significantly reduced this problem by introducing multi-stage eigenanalysis-based extensions of our approach, where the term “multi-stage” refers to the fact these methods successively perform several eigendecompositions. We explained in detail how two-stage methods (especially EQPT2 and EQPT3) operate, and we showed how to extend them to an arbitrary number of stages in order to minimize their sensitivity to QST errors (especially with the dichotomic EQPT5 method).
@@
Appendix A Mathematical definitions and properties
@@
See [43] p. 157 (e). See also [43] p. 101 for unitarily diagonalizable matrices (every normal matrix is unitarily diagonalizable) and the spectral theorem for normal matrices. In particular, this implies the spectral theorem for Hermitian matrices: see [43] p. 104.
@@
where $v_{k}$ is the $k$th column of $P$ and $\lambda_{k}$ is the $k$th element on the diagonal of $D$. Eq. (42) shows that $v_{k}$ is an eigenvector of $A$ and that the associated eigenvalue is $\lambda_{k}$.
@@
A best least-squares approximation of a given matrix $A\in M_{n}$ by a unitary matrix $U_{A}\in M_{n}$ is given by $U_{A}=VW^{{\dagger}}$, where $A=V\Sigma W^{{\dagger}}$ is a singular value decomposition of $A$.
@@
See [43] p. 431.
Appendix B A variant for the first part of single-stage QPT methods
We here again consider the first part of single-stage QPT methods. As in the variant that we proposed in Section II.1 for that first part, we here use a mixed input state, represented by the density matrix $\rho_{1}$. However, for the sake of generality, this matrix is here not any more constrained to be diagonal. It is still Hermitian and therefore unitarily diagonalizable. In the version considered here, as in Section II.1, we request all its eigenvalues to be different. That input density matrix $\rho_{1}$ is here requested to be known (i.e. fixed a priori or estimated from copies of that state, so that this variant is supervised or semi-supervised).
Using the same properties of eigendecomposition as in Section II.1, we consider all unitarily diagonalized forms of $\rho_{1}$ that have the following properties:
@@ Appendix B A variant for the first part of single-stage QPT methods
Starting from a given $\rho_{1}$, the first step of the algorithm proposed in the present section consists of determining one (arbitrary) of these decompositions (43) meeting the above-defined conditions. This is performed by 1) running an eigendecomposition algorithm, then 2) reordering the values in the matrix obtained for $\rho_{I1D}$ if they are not directly provided in decreasing order by the considered eigendecomposition algorithm, 3) reordering the columns of the matrix obtained for $U_{I1}$ accordingly and 4) dividing these columns by their norms if the considered practical eigendecomposition algorithm is not guaranteed to yield unit-norm vectors. Thus, we especially get a known value $U_{I1}$ (which contains indeterminacies consisting of a phase factor separately for each of its columns). $U_{I1}$ is used in the second part of this version of our algorithm: see Appendix D.
@@
Formally, (44) adresses exactly the same problem as in Section II.1 (see (1) with a diagonal $\rho_{1}$), but now with a diagonal input density matrix $\rho_{I1D}$ and a process $U_{g}$ (that is also unitary). We can therefore apply the algorithm and results of Section II.1 to this reformulated problem, i.e., briefly, we diagonalize $\widehat{\rho}_{4}$ and we then reorder the eigenvalues (in the same order as in $\rho_{I1D}$) and associated normalized eigenvectors. This guarantees that the algorithm of Section II.1 here provides an estimate $\widehat{U}_{2}$ that restores $U_{g}$ up to a phase factor for each of its columns, with (7) here replaced by
@@
As opposed to the basic version of the first step defined in Section II.1, the estimate $\widehat{U}_{2}$ here contains the effect of $U_{I1}$ in addition. The second part of the method must then be modified accordingly, as compared with its basic version defined in Section II.2. A solution to this problem is provided in Appendix D.
Appendix C A variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods that include a first part as defined in Section II.1, we here again consider the associated second part, that aims at estimating the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, by using a second input state of the considered process. The method that we proposed in Section II.2 for that second part uses a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. However, a mixed input state having a density matrix $\rho_{5}$ may be used instead, as will now be shown. As $\lvert\Psi_{1}\rangle$ in Section II.2, the input density matrix $\rho_{5}$ is here supposed to be initially known (supervised case) or estimated by using part of the available copies of that state (semi-supervised case).
@@ Appendix C A variant for the second part of single-stage QPT methods
Then, knowing $\widehat{\rho}_{8}$ and $\widehat{U}_{2}$ obtained in the first part of our algorithm (see Section II.1), the next step of our algorithm here consists of combining them to form
@@
and the state $\lvert\Psi_{3}\rangle$ considered in Section II.1 corresponds to $\rho_{9}=\lvert\Psi_{3}\rangle\langle\Psi_{3}\lvert$. Combining the latter expression with (12) and using (50), the resulting expression of $\rho_{9}$ is fully compatible with its extension (49) to an arbitrary mixed input state considered here.
@@
and $\rho_{9}$ does not yield any information about the phase factors in $D_{\phi}$. This is not surprising because all the information that can be extracted with a diagonal input density matrix was already extracted by using such a matrix in the first part of our algorithm (see Section II.1). We therefore consider a non-diagonal matrix $\rho_{5}$ hereafter.
@@
Appendix D Another variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods, we here again consider the case when the first part is defined as in Appendix B. We present an associated solution for the second part of the method, using a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. $\lvert\Psi_{1}\rangle$ is requested to be known, i.e. fixed a priori (supervised case) or estimated from a subset of its copies (semi-supervised case).
Starting with the same approach as in Section II.2, this here again yields the output ket estimate (11). Our algorithm then computes $\lvert\Psi_{3}\rangle$ defined by (12), as in Section II.2, but now with $\widehat{U}_{2}$ that meets (47). Combining (47) with (12) and (11) and considering that $U$ is unitary (hence $U^{{\dagger}}U=I$) , we get
@@ Appendix D Another variant for the second part of single-stage QPT methods
which is a known quantity. Comparing (60) with (13) shows that $\lvert\Psi_{4}\rangle$ here replaces the ket $\lvert\Psi_{1}\rangle$ used in the method of Section II.2 (which is coherent with the fact that the latter method is a reduced form of the present one, for the case when $\rho_{1}$ is diagonal and hence $U_{I1}$ is the identity matrix). The remainder of the present method is then derived by modifying the end of the method of Section II.2 accordingly: here using
@@
it may be shown in the same way as in Section II.2 that we here again get (17), i.e. $\widehat{U}_{5}$ here again succeeds in restoring $U$, up to a global phase factor (and again up to estimation errors).
Appendix E Variants of two-stage QPT methods
@@
Since the columns of $\widehat{U}_{4}$ are estimated independently, one may fear that they are not exactly orthogonal, due to estimation errors. Then, $\widehat{U}_{4}$ and hence $\widehat{U}_{5}$ are not unitary, although they should be because they aim at estimating $U$, that is constrained to be unitary. A first solution to this problem consists of postprocessing $\widehat{U}_{4}$, so as to determine a best least-squares unitary approximation of $\widehat{U}_{4}$. An algorithm that may be used to this end is defined by Theorem 4. The resulting modified QPT method is called EQPT3 hereafter. A slightly different version consists of finding a unitary approximation of $\widehat{U}_{5}$ instead of $\widehat{U}_{4}$. The resulting QPT method is called EQPT4.
Appendix F Relevance of the proposed multi-stage methods and pseudo-code
@@
The successive subspace intersections may be considered in such an order that the sets of indices $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$ correspond to the representation in base 2 of the integers in increasing order. Then, the columns of $U$ are obtained with indices in increasing order. This is exactly what is obtained in the successive rows of Table 1.

### v2

@@ VI Conclusion and future works
The above approach may however be expected to yield limited performance when considering high-dimensional state spaces, due to the estimation errors of the Quantum State Tomography (QST) that it uses as a building block. We very significantly reduced this problem by introducing multi-stage eigenanalysis-based extensions of our approach, where the term “multi-stage” refers to the fact these methods successively perform several eigendecompositions. We explained in detail how two-stage methods (especially EQPT2 and EQPT3) operate, and then how to extend them to an arbitrary number of stages in order to minimize their sensitivity to QST errors (especially with the dichotomic EQPT5 method).
@@
Appendix A Connections between our QPT methods and classical signal processing
As stated in Section I, the type of QPT approaches proposed in this paper has connections with works that were previously reported in the classical domain. These connections may be defined as follows. QPT may be considered as the quantum version of so-called (classical) system identification [1]. The latter classical problem consists of estimating the parameter values of a transform from the input to the output of a system. That system may have multiple inputs, so that it then performs a combination of them. In the classical framework, such a combination is called a mixture. The classical problem then considered in the literature is therefore also referred to as mixture identification and it especially includes blind mixture identification, or BMI [59, 58, 57, 60]. This terminology should be used with care when also considering the quantum domain, because the above classical “mixtures” have nothing to do with the “statistical mixtures”, i.e. “mixed states”, faced in the quantum domain and used further in this paper. Only considering previous works in the classical domain at this stage of our discussion, the above-mentioned BMI has a close relationship with source separation, especially including blind source separation, or BSS [76, 77, 58, 57, 50, 78, 53]. BSS does not focus on identifying the transform performed by the above, i.e. direct, system itself, but mainly on deriving an estimate of the inverse of that transform, so as to apply it to the outputs of the above direct system and to thus restore the unknown source signals that form the input of the direct system. This relationship between BMI and BSS even appears in the name of the well-known SOBI method [79], which stands for “Second-Order Blind Identification”, whereas it is claimed to be a BSS method. The term “Second-Order” here refers to the fact that this method applies to random signals and uses the statistical properties of the measured signals, more precisely their usual second-order statistical parameters that consist of their covariance matrices involving different time delays between signals.
So far, we especially discussed how SOBI-like methods and the methods proposed in the present paper are connected in terms of the nature of data processing problems that they address. These two types of methods also have relationships in terms of the approaches that they use to solve the considered problems. More precisely, in classical BSS/BMI, when the transform performed by the direct system is linear (and memoryless) and its identification is performed by using the above-mentioned second-order statistics of the output signals of the direct system, an equation similar to (1) is obtained [80, 79, 60, 81, 82]. That classical counterpart of the structure of (1) was already exploited to perform classical BSS/BMI by means of several methods based on eigendecomposition. The real-valued version of the basic structure of (1) was first used in [80, 60] together with a required modified form of this equation (a related approach also appears in [82]). It was then extended in [79], by combining it with more than one modified form of this equation, to obtain a more robust method. Another version [81] aims at extending that type of approach from the standard case of stationary time-domain source signals to nonstationary time-domain sources. That version eventually uses complex-valued signals (whereas the mixing transform remains real-valued) because one moves to the Fourier domain.
Appendix B Creating a mixed state that has a diagonal density matrix
The algorithm proposed in Section II-A uses an input density matrix $\rho_{1}$ that is diagonal. It should be noted that, when one considers that one is able to create a pure state, then, creating the type of state considered here is not a problem, as will now be shown. We here consider the approach initially used by von Neumann in [83] to define a statistical mixture: such a state is expressed as being composed of a set of pure states, each of them being associated with a probability of occurrence. Then, to create experimental data corresponding to the state defined by the diagonal density matrix $\rho_{1}$, one just has to repeatedly, randomly, select one of the pure states that form the considered basis of the state space, moreover selecting the $k$th basis state with a probability that is equal to the $k$th value on the diagonal of $\rho_{1}$. This simplicity of implementation is worth noting because, in constrast, the authors of [44, 3] consider that using a mixed input state is too complex (maybe because they have non-diagonal density matrices in mind), so that they choose to restrict themselves to pure states, which has the drawback of requesting many such multiqubit input states, as opposed to only two multiqubit input states in the complete approach that we here start to describe (namely the mixed state with a density matrix $\rho_{1}$ considered here, and then the pure state $\lvert\Psi_{1}\rangle$ in Section II-B). Beyond their relationship based on considering (1), our approach and the one in [44] are therefore quite different.
Appendix C Preprocessing steps of the single-stage QPT algorithm
The first part of our single-stage algorithm, described in Section II-A, is based on an estimated density matrix $\widehat{\rho}_{2}$ provided by a QST algorithm. If this estimate $\widehat{\rho}_{2}$ is not guaranteed to be Hermitian due to the estimation errors of the considered “external” (i.e. external to our QPT) QST algorithm, a first recommended preprocessing step of our method consists of deriving a Hermitian estimate $\widehat{\rho}_{3}$ from $\widehat{\rho}_{2}$. This may be achieved by computing the Hermitian part of $\widehat{\rho}_{2}$, that is (see [48] p. 109 or [50] p. 343 for its real-valued counterpart):
Similarly, the second part of our single-stage algorithm, described in Section II-B, uses an estimated ket $\lvert\widehat{\Psi}_{2}\rangle$, provided by a QST algorithm. If the considered QST algorithm does not guarantee that the norm of $\lvert\widehat{\Psi}_{2}\rangle$ is equal to 1, a recommended pre-processing step of our QPT method consists of dividing $\lvert\widehat{\Psi}_{2}\rangle$ by its norm.
Appendix D Preferred input density matrix for the first part of the single-stage QPT algorithm
In the first step of our first method defined in Section II-A, we preferably use the value of $\rho_{1}$ defined as follows. The values on its diagonal are the eigenvalues of the considered density operator $\rho_{2}$, they are real and non-negative and their sum is equal to one (see e.g. [1] p. 101). As shown in Section II-A, the proposed method assumes that these eigenvalues are different and ordered, and it is based on reordering the eigenvalues in the eigendecomposition of the estimated output density matrix and on permuting the eigenvectors accordingly. Therefore, if using quite close actual eigenvalues when choosing $\rho_{1}$, one may fear that the corresponding estimated eigenvalues obtained for the output density matrix be permuted due to estimation errors, so that the corresponding eigenvectors may then be permuted accordingly. This is a major problem, because the intermediate estimate $\widehat{U}_{2}$ related to the considered process, obtained at the output of this first step of our QPT method, is primarily based on these eigenvectors. In this paper, we handle this problem as follows: first, for the method imposed in Section II-A, we select the value of $\rho_{1}$ in order to limit that problem to the greatest extent that is possible within the frame of that method; then, in Sections III and IV, we propose more advanced methods that allow one to further reduce that problem. So, for the method considered here, we choose eigenvalues that are uniformly distributed (i.e. multiples of the same step, with zero excluded), so as to maximize the gap between any two adjacent values. Moreover constraining them to be positive, with a sum equal to one and in decreasing order as explained above, it is easily shown that one obtains the following values for the diagonal elements $\rho_{1,kk}$ of $\rho_{1}$:
Appendix E Preferred input ket for the second part of the single-stage QPT algorithm
In the second step of first method defined in Section II-B, we select the input ket $\lvert\Psi_{1}\rangle$ as follows. Based on the condition on nonzero components provided in Section II-B for $\lvert\Psi_{1}\rangle$, we preferably select a state $\lvert\Psi_{1}\rangle$ with all component moduli as high as possible. Considering positive real-valued components and taking into account that $\lvert\Psi_{1}\rangle$ has unit norm, this yields a single solution: all components of $\lvert\Psi_{1}\rangle$ in the considered basis are equal to $1/\sqrt{d}$.
Appendix F Mathematical definitions and properties
@@
See [48] p. 157 (e). See also [48] p. 101 for unitarily diagonalizable matrices (every normal matrix is unitarily diagonalizable) and the spectral theorem for normal matrices. In particular, this implies the spectral theorem for Hermitian matrices: see [48] p. 104.∎
@@
where $v_{k}$ is the $k$th column of $P$ and $\lambda_{k}$ is the $k$th element on the diagonal of $D$. Eq. (44) shows that $v_{k}$ is an eigenvector of $A$ and that the associated eigenvalue is $\lambda_{k}$.∎
@@
A best least-squares approximation of a given matrix $A\in M_{n}$ by a unitary matrix $U_{A}\in M_{n}$ is given by $U_{A}=VW^{{\dagger}}$, where $A=V\Sigma W^{{\dagger}}$ is a singular value decomposition (SVD) of $A$.
@@
See [48] p. 431.∎
Appendix G A variant for the first part of single-stage QPT methods
We here again consider the first part of single-stage QPT methods. As in the variant that we proposed in Section II-A for that first part, we here use a mixed input state, represented by the density matrix $\rho_{1}$. However, for the sake of generality, this matrix is here not any more constrained to be diagonal. It is still Hermitian and therefore unitarily diagonalizable. In the version considered here, as in Section II-A, we request all its eigenvalues to be different. That input density matrix $\rho_{1}$ is here requested to be known (i.e. fixed a priori or estimated from copies of that state, so that this variant is supervised or semi-supervised).
Using the same properties of eigendecomposition as in Section II-A, we consider all unitarily diagonalized forms of $\rho_{1}$ that have the following properties:
@@ Appendix G A variant for the first part of single-stage QPT methods
Starting from a given $\rho_{1}$, the first step of the algorithm proposed in the present section consists of determining one (arbitrary) of these decompositions (45) meeting the above-defined conditions. This is performed by 1) running an eigendecomposition algorithm, then 2) reordering the values in the matrix obtained for $\rho_{I1D}$ if they are not directly provided in decreasing order by the considered eigendecomposition algorithm, 3) reordering the columns of the matrix obtained for $U_{I1}$ accordingly and 4) dividing these columns by their norms if the considered practical eigendecomposition algorithm is not guaranteed to yield unit-norm vectors. Thus, we especially get a known value $U_{I1}$ (which contains indeterminacies consisting of a phase factor separately for each of its columns). $U_{I1}$ is used in the second part of this version of our algorithm: see Appendix I.
@@
Formally, (46) adresses exactly the same problem as in Section II-A (see (1) with a diagonal $\rho_{1}$), but now with a diagonal input density matrix $\rho_{I1D}$ and a process $U_{g}$ (that is also unitary). We can therefore apply the algorithm and results of Section II-A to this reformulated problem, i.e., briefly, we diagonalize $\widehat{\rho}_{4}$ and we then reorder the eigenvalues (in the same order as in $\rho_{I1D}$) and associated normalized eigenvectors. This guarantees that the algorithm of Section II-A here provides an estimate $\widehat{U}_{2}$ that restores $U_{g}$ up to a phase factor for each of its columns, with (5) here replaced by
@@
As opposed to the basic version of the first step defined in Section II-A, the estimate $\widehat{U}_{2}$ here contains the effect of $U_{I1}$ in addition. The second part of the method must then be modified accordingly, as compared with its basic version defined in Section II-B. A solution to this problem is provided in Appendix I.
Appendix H A variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods that include a first part as defined in Section II-A, we here again consider the associated second part, that aims at estimating the phase factors $e^{i\phi_{1}},\dots,e^{i\phi_{d}}$, by using a second input state of the considered process. The method that we proposed in Section II-B for that second part uses a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. However, a mixed input state having a density matrix $\rho_{5}$ may be used instead, as will now be shown. As $\lvert\Psi_{1}\rangle$ in Section II-B, the input density matrix $\rho_{5}$ is here supposed to be initially known (supervised case) or estimated by using part of the available copies of that state (semi-supervised case).
@@ Appendix H A variant for the second part of single-stage QPT methods
Then, knowing $\widehat{\rho}_{8}$ and $\widehat{U}_{2}$ obtained in the first part of our algorithm (see Section II-A), the next step of our algorithm here consists of combining them to form
@@
and the state $\lvert\Psi_{3}\rangle$ considered in Section II-A corresponds to $\rho_{9}=\lvert\Psi_{3}\rangle\langle\Psi_{3}\lvert$. Combining the latter expression with (9) and using (52), the resulting expression of $\rho_{9}$ is fully compatible with its extension (51) to an arbitrary mixed input state considered here.
@@
and $\rho_{9}$ does not yield any information about the phase factors in $D_{\phi}$. This is not surprising because all the information that can be extracted with a diagonal input density matrix was already extracted by using such a matrix in the first part of our algorithm (see Section II-A). We therefore consider a non-diagonal matrix $\rho_{5}$ hereafter.
@@
Appendix I Another variant for the second part of single-stage QPT methods
In the framework of single-stage QPT methods, we here again consider the case when the first part is defined as in Appendix G. We present an associated solution for the second part of the method, using a pure input state defined by a ket $\lvert\Psi_{1}\rangle$. $\lvert\Psi_{1}\rangle$ is requested to be known, i.e. fixed a priori (supervised case) or estimated from a subset of its copies (semi-supervised case).
Starting with the same approach as in Section II-B, this here again yields the output ket estimate (8). Our algorithm then computes $\lvert\Psi_{3}\rangle$ defined by (9), as in Section II-B, but now with $\widehat{U}_{2}$ that meets (49). Combining (49) with (9) and (8) and considering that $U$ is unitary (hence $U^{{\dagger}}U=I$) , we get
@@ Appendix I Another variant for the second part of single-stage QPT methods
which is a known quantity. Comparing (62) with (10) shows that $\lvert\Psi_{4}\rangle$ here replaces the ket $\lvert\Psi_{1}\rangle$ used in the method of Section II-B (which is coherent with the fact that the latter method is a reduced form of the present one, for the case when $\rho_{1}$ is diagonal and hence $U_{I1}$ is the identity matrix). The remainder of the present method is then derived by modifying the end of the method of Section II-B accordingly: here using
@@
it may be shown in the same way as in Section II-B that we here again get (14), i.e. $\widehat{U}_{5}$ here again succeeds in restoring $U$, up to a global phase factor (and again up to estimation errors).
Appendix J Variants of two-stage QPT methods
@@
Since the columns of $\widehat{U}_{4}$ are estimated independently, one may fear that they are not exactly orthogonal, due to estimation errors. Then, $\widehat{U}_{4}$ and hence $\widehat{U}_{5}$ are not unitary, although they should be because they aim at estimating $U$, that is constrained to be unitary. A first solution to this problem consists of postprocessing $\widehat{U}_{4}$, so as to determine a best least-squares unitary approximation of $\widehat{U}_{4}$. An algorithm that may be used to this end is defined by Theorem 4. The resulting modified QPT method is called EQPT3 hereafter. Its pseudo-code is provided in Algorithm 3. A slightly different version consists of finding a unitary approximation of $\widehat{U}_{5}$ instead of $\widehat{U}_{4}$. The resulting QPT method is called EQPT4. Its pseudo-code is provided in Algorithm 4.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       /* Exploit $\widehat{\rho}_{8}$: */
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
       for $m_{1}=1$ to $d_{2}$ do
             for $m_{2}=1$ to $d_{2}$ do
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
                   /* If the CCA tool does not provide unit-norm vectors, then do it: */
                   Divide above column of $\widehat{U}_{4}$ by its norm;
             end for
       end for
      Find decomposition of $\widehat{U}_{4}$ as $\widehat{U}_{4}=V\Sigma W^{{\dagger}}$ with SVD;
       $\widehat{U}_{4}{}_{u}=VW^{{\dagger}}$;
       /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}{}_{u}\hskip 0.28453pt^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{4}{}_{u}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
end
Algorithm 3 EQPT3: third Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by QST and obtained for input density matrix (17) with the diagonal values defined in (19)). b) Estimate $\widehat{\rho}_{8}$ of output density matrix $\rho_{6}$ (provided by QST and obtained for input density matrix (21) with the diagonal values defined in (19)). c) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
Output : Estimate $\widehat{U}_{5}$ of quantum process matrix $U$.
begin
       /* Exploit $\widehat{\rho}_{4}$: */
       /* Exploit $\widehat{\rho}_{8}$: */
       /* Combine $\widehat{U}_{2}$ and $\widehat{U}_{3}$: */
       for $m_{1}=1$ to $d_{2}$ do
             for $m_{2}=1$ to $d_{2}$ do
                   Set column of $\widehat{U}_{4}$ with index equal to $((m_{1}-1)d_{1}+m_{2})$: Set it to first vector provided by CCA of 1) the set of columns of $\widehat{U}_{2}$ with indices $((m_{1}-1)d_{1}+n_{1})$ with $n_{1}\in\{1,\dots,d_{1}\}$ and 2) the set of columns of $\widehat{U}_{3}$ with indices $((m_{2}-1)d_{1}+n_{2})$ with $n_{2}\in\{1,\dots,d_{1}\}$;
                   /* If the CCA tool does not provide unit-norm vectors, then do it: */
                   Divide above column of $\widehat{U}_{4}$ by its norm;
             end for
       end for
      /* Exploit $\lvert\widehat{\Psi}_{2}\rangle$: */
       $\lvert\Psi_{3}\rangle=\widehat{U}_{4}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}{}_{p}\hskip 0.28453pt=\widehat{U}_{4}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
       Find decomposition of $\widehat{U}_{5}{}_{p}$ as $\widehat{U}_{5}{}_{p}=V\Sigma W^{{\dagger}}$ with SVD;
       $\widehat{U}_{5}=VW^{{\dagger}}$;
end
Algorithm 4 EQPT4: fourth Eigenanalysis-based QPT algorithm (composed of two stages). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see text.
Appendix K Additional information for the proposed multi-stage methods
K-A Relevance of the methods
@@
The successive subspace intersections may be considered in such an order that the sets of indices $\{\mu_{0},\dots,\mu_{{b_{\ell m}}}\}$ correspond to the representation in base 2 of the integers in increasing order. Then, the columns of $U$ are obtained with indices in increasing order. This is exactly what is obtained in the successive rows of Table II in the example provided below.
K-B Example
We here illustrate the above property and thus the proposed complete multi-stage methods by considering the case $d=8$, $d_{2}=2$ and hence $d_{1}=4$ and $(\log_{2}d_{1}+1)=3$ stages. During the first stage, that corresponds to ${b_{\ell}}=0$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{0,m_{0}}$, respectively with $m_{0}=1$ and $m_{0}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,3,4\}$ and $\{5,6,7,8\}$. Then, during the second stage, that corresponds to ${b_{\ell}}=1$, the eigendecomposition allows one to identify two subspaces ${\cal S}_{1,m_{1}}$, respectively with $m_{1}=1$ and $m_{1}=2$, and that respectively correspond to the following sets of 4 indices of columns of $U$: $\{1,2,5,6\}$ and $\{3,4,7,8\}$. Finally, the third stage, that corresponds to ${b_{\ell}}=2$, yields the following sets of 4 indices of columns of $U$: $\{1,3,5,7\}$ and $\{2,4,6,8\}$. We then derive the corresponding subspace intersections, first considering ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}$ as intermediate quantities, and finally deriving ${\cal S}_{0,m_{0}}\cap{\cal S}_{1,m_{1}}\cap{\cal S}_{2,m_{2}}$. All these intersections are listed in Table II. This proves that the proposed approach yields 8 intersections that are each restricted to one column of $U$ and that this set of intersections yields all the individual columns of $U$, as required.
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{1,2\}$$\{1\}$
$\{1,2,3,4\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{1,2\}$$\{2\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{3,4\}$$\{3\}$
$\{1,2,3,4\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{3,4\}$$\{4\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{1,3,5,7\}$$\{5,6\}$$\{5\}$
$\{5,6,7,8\}$$\{1,2,5,6\}$$\{2,4,6,8\}$$\{5,6\}$$\{6\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{1,3,5,7\}$$\{7,8\}$$\{7\}$
$\{5,6,7,8\}$$\{3,4,7,8\}$$\{2,4,6,8\}$$\{7,8\}$$\{8\}$
TABLE II: Multi-stage QPT method. Each column of the table contains the indices of the column vectors that correspond to the considered subspace or subspace intersection defined in the first row of the table.
K-C Pseudo-code